# Three Dimensional Topological Field Theories and Nahm Sum Formulas

### v1

@@ 4.2 Unitary TFTs
${\cal I}^{K}_{\rm sci}(q,\vec{\mu},\vec{\zeta}=\vec{1})=1\textrm{for all $\vec{\mu}\in\mathfrak{M}[K,\{{\cal O}_{I}\}]$\ ,}$ (57)
which implies that the only local operator in the unitary TFT is the identity operator. Similar to the non-unitary case, the UV gauge theory has a flavor symmetry whose rank is equal to the dimension of $\mathfrak{M}[K,\{{\cal O}_{I}\}]$. For some examples the index is $1$ for all $\vec{\mu}\in\mathfrak{M}[K,\{{\cal O}_{I}\}]$, even though its dimension is non-zero. In this case, it is natural to expect that the UV flavor symmetry decouples in the IR. As a consequence, the superconformal index and the three-sphere free energy defined in (27) do not depend on the mixing parameter $\vec{\mu}\in\mathfrak{M}[K,\{{\cal O}_{I}\}]$.
For a UV supersymmetric Wilson loop $W_{\vec{Q}}$ to be a simple object in the IR TFT, it must satisfy conditions similar to those in (33):
$\begin{split}&(i)\;\;\langle W^{+}_{\vec{Q}}\rangle_{\rm sci}(q,\vec{\mu},\vec{\zeta}=1)=0\textrm{ or }\pm q^{\mathbb{Z}/2}\;,\\ &(ii)\;\;\langle W^{+}_{\vec{Q}}W^{-}_{\vec{Q}}\rangle_{\rm sci}(q,\vec{\mu},\vec{\zeta}=1)=1\;,\end{split}$ (58)
@@
If ${\cal T}[K,\{{\cal O}_{I}\}]$, with positive definite $K$, has a mass gap and a UV loop $W_{\vec{Q}}$ flows to a simple object in the IR TFT, the half-index
@@ 4.2.2 $r=2$
$\begin{split}&U(1)_{a-1}\textrm{ coupled to $\Phi_{1}$(of charge $+1$) and $\Phi_{2}$ (of charge $-1$) with superpotential ${\cal W}=\Phi_{1}\Phi_{2}$}\\ &\simeq U(1)_{a-1}\textrm{ Chern-Simons theory.}\end{split}$
The first line follows from the basic mirror symmetry reviewed in Section 41. Integrating out the two chiral fields, we obtain the pure CS theory in the infrared. As discussed in Dimofte:2017tpi , the deformed Dirichlet boundary condition for $\Phi_{1}$ breaks the supersymmetry spontaneously, unless its boundary value is at the critical point of the superpotential.
@@
The half-index computation suggests that the only boundary operator that survives in this case is the identity operator. It is interesting to note that we also have  [^13] It is tempting to conjecture that the deformed Dirichlet boundary condition flows to that corresponds to “closing the puncture” of $D^{2}\times S^{1}$ in the IR, which produces the partition function on $S^{2}\times S^{1}$.
@@ 4.3 Comparision with Zagier’s result
Let us now compare our results with those in Zagier:2007knq . The half-indices $I_{A}[W_{\vec{Q}}]$ in (39) and $I_{\rm half}^{W_{\vec{Q}}}(q,\vec{\mu})$ in (59) correspond to the Nahm sum formula $\chi_{(A,B,C)}$ in (3) with the identification
@@
provided that $B\in\mathbb{Z}^{r}$. We find that all the triplets $(A,B,C)$ in Zagier’s list, summarized in Table 1 (with $B\in\mathbb{Z}^{r}$), appear in our classification except for two cases.
@@
In this case ${\cal T}[K,\emptyset]$ has only one primitive 1/2 BPS monpole operator $V_{(0,1,1)}$. After the superpotential deformation, the gauge theory ${\cal T}[K,V_{(0,1,1)}]$ flows to an ${\cal N}=4$ SCFTs in the IR. The UV gauge theory has two $U(1)$ topological symmetries but the computation of superconformal index strongly suggests that one linear combintations of them acts trivially in the IR. If this happens, the example does not show up in our search due to our simplifying assumption (a) in section 3.2.
@@
Here $\alpha\in\mathbb{R}$ parametrize the mixing of R-symmetry with the decoupled $U(1)$ symmetry. One can check that the superconformal index is independent of $\alpha$. The half-index computation at $\alpha=0$ becomes
$\begin{split}I_{\rm half}(q,\nu=-1,\eta^{-1})&=~{}\sum_{m_{1},m_{2},m_{3}\geq 0}\frac{q^{\frac{1}{2}m^{t}Km}\eta^{m_{2}-m_{3}}}{(q)_{m_{1}}(q)_{m_{2}}(q)_{m_{3}}}\\ &=\left(I_{\rm half}(q,\nu=1,\eta)[{\cal T}_{\text{\text@underline{1-1}}}]\right)\times\left(\frac{1}{(q)_{\infty}}\sum_{m\in\mathbb{Z}}q^{am^{2}/2}\eta^{-m}\right)\,.\end{split}$ (69)
@@
which suggests that ${\cal T}[K,\{V_{(0,1,1)}\}]$ is a mirror dual of ${\cal T}_{\text{\text@underline{1-1}}}\otimes U(1)_{a}$.
@@
in Zagier’s list. The Nahm sum formula with (71) for either choice of $B$ is
@@
This is a modular function but we do not know if it can be identified with a character of any rational VOA. [^15] As coefficients of $\chi^{U(1)_{2}}_{1}(q)$ are all even, the coefficients of $q$-series are all integer-valued It can be reproduced from the half index of U3-1 with $a=2$ at $\vec{\mu}=\vec{0}$ decorated by a supersymmetric Wilson loop of charge $\vec{Q}=-(1,1,0)$. However, when computing the superconformal index:
@@ 5 Discussion
The gauge theory ${\cal T}[K,\{\mathcal{O}_{I}\}_{I=1}^{r-1}]$ should be understood as a sequence of renormalization group (RG) flows:
${\cal T}[K,\emptyset]$ $\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{1}\;}{\cal T}[K,\{\mathcal{O}_{1}\}]\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{2}\;}\ldots\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{r-1}\;}{\cal T}[K,\{\mathcal{O}_{I}\}_{I=1}^{r-1}]\;.$ (74)
For these RG flows to make sense, the superpotential at each step must be relevant:
@@
Here, $R^{(I-1)}_{0}=R_{*}+\vec{\mu}^{(I-1)}_{0}\cdot\vec{T}$ represents the superconformal R-charge of ${\cal T}[K,\{\mathcal{O}_{A}\}_{A=1}^{I-1}]$, which can be determined by extremizing $F(\vec{\mu})$ over $\vec{\mu}\in\mathfrak{M}[K,\{\mathcal{O}_{1},\ldots,\mathcal{O}_{I-1}\}]$. Thus, in order for the fully deformed theory, ${\cal T}[K,\{\mathcal{O}_{I}\}_{I=1}^{r-1}]$, to be sensible, there should be a proper sequence of 1/2-BPS operators, $\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{r-1}$, which ensures that these relevance conditions are satisfied.
Verifying the relevance of the superpotential is a challenging task and will be addressed in future work. Instead, we will list inconsistencies which indirectly demonstrate that certain examples in our classification may violate this condition:
  • In the cases of 3-14-(2) and 3-20, the half indices $I_{A}[W_{\vec{Q}}]$ contain a factor $\chi_{F}^{-1}$. This does not correspond to any known RCFT character in the literature.
  • For the 3-14-(2) case, the A and B-twisted partition functions on $S^{2}\times S^{1}$ are not equal to $1$.
  • In the 3-20 case, the half-index $I_{A}[W_{(1,1,0)}]$ is expressed as a sum of two characters (modulo a $\chi_{F}^{-1}$), even though $W_{(1,1,0)}\in W^{\rm sim}_{A}$.
  • Notably, both 3-14-(2) and 3-20 involve a non-primitive monopole operator: $V_{(-2,0,4)}$ in 3-14-(2) and $V_{(-2,2,2)}$ in 3-20, which we suspect to be irrelevant.
@@
In section 3.3, we have scanned UV Wilson loops which flow to simple objects in the IR TFT. Half-indices with Wilson loops insertions matched to RCFT characters of corresponding primaries. There are some exceptional cases where we could not find the UV Wilson loop corresponding to an IR simple object:
[…]
In 3-22, we could not find the UV Wilson loop corresponding to $\chi^{SM(2,12)}_{(1,3)}$. Instead, we find
@@
[…]
@@
$q^{2}I_{A}[W_{(-2,2,1)}]=q^{4}I_{A}[W_{(-4,3,1)}]=q^{-3/28}\chi^{W_{3}(3,7)}_{(3,3,1)}.$ (81)
However, they are not simple:
$\langle W^{-}_{(-2,2,1)}W^{+}_{(-2,2,1)}\rangle_{\rm sci}=\langle W^{-}_{(-4,3,1)}W^{+}_{(-4,3,1)}\rangle_{\rm sci}=3.$ (82)
This suggests that the Wilson loops can be expressed as linear combinations of three IR simple objects. Based on our computations, we propose the following expressions:
@@
Note that there are two primaries with a conformal dimension of $-3/7$ in the $W_{3}$ minimal model, and the corresponding bulk simple objects are denoted by $\mathbf{L}_{-\frac{3}{7}}$ and $\widetilde{\mathbf{L}}_{-\frac{3}{7}}$. They are related to each other by charge conjugation and they share the same conformal character. The above identification is also consistent with $\langle W^{-}_{(-2,2,1)}W^{+}_{(-4,3,1)}\rangle_{\rm sci}=-q^{-2}$.
@@ A.1 General structure of partition functions on $\mathcal{M}_{g,p}$
The effective twisted superpotential $\mathcal{W}(u,v)$ is broken down into two pieces, one originating from the Chern-Simons terms of the theory and one originating from the one-loop contribution of the $\Phi_{i}$ [^17] Importantly this is all done in the “$U(1)_{-\frac{1}{2}}$ quantization” which is arises due to the parity anomaly that exists in 3D gauge theories as discussed in Closset:2018ghr .:
@@
where the A subscript indicates we are utilizing the A-twist background on $S^{2}$ as described in Closset:2017zgf . The flavor flux operator $\Pi_{i}(u,v)^{\mathfrak{n}_{i}}$ can also be inserted to compute the partition function in the presence of a nontrivial flavor flux $\mathfrak{n}_{i}$ threaded through the base space $\Sigma_{g}$. Lastly, the gauge flux operator, $\Pi_{a}(u,v)$ is a rational function in the $x_{a}$ which defines the Bethe vacua [^18] For non-abelian gauge group $G$ the action of the Weyl group of $G$ on ${\cal S}_{BE}$ must be considered. as solutions to the system of polynomial equations:
@@
$\mathcal{Z}_{\mathcal{M}_{g,p}}(v;v_{R},\mathfrak{n}_{i})=\sum_{\hat{u}\in\mathcal{S}_{BE}(v;v_{R})}\mathcal{H}(\hat{u},v)^{g-1}\mathcal{F}(\hat{u},v)^{p}\prod_{i}\Pi_{i}(\hat{u},v)^{\mathfrak{n}_{i}}.$ (94)
@@ A.2.1 $U(1)_{R}$ Holonomy and Flux in the 3D A-model
Recall that for a field with R-charge $r$ the Dirac quantization condition is $r\mathfrak{n}_{R}\in\mathbb{Z}$ and thus the A-model background generically enforces $r\in\mathbb{Z}$. However, examining the conditions (96), we see that when $p\neq 0$, $g-1=0\ (\text{mod}\ p)$, and $v_{R}$ is integral [^19] One can find conditions for $v_{R}$ half-integral, namely $g-1=\frac{p}{2}\ (\text{mod}\ p)$, but this will not be important here., a specific representative $v_{R}$ can be chosen such that $\mathfrak{n}_{R}=0$ in which case we are free to allow $r\in\mathbb{R}$. This is crucial for F-maximization where the R-charge must be allowed to take on arbitrary real values.
@@ A.2.2 F-maximization in the 3D A-model
Using this we see how the functions of section A.1 change: ${\cal W}_{\text{CS}}$ gains a term of the form $v_{R}\vec{u}\cdot\vec{\mu}$ and the flux operators $\Pi_{i}$ appearing in (94) have their exponent shifted by $\mu_{i}\mathfrak{n}_{R}$. The latter fact indicates that if we are working in a background with no flavor flux threaded through $\Sigma_{g}$ then the R/flavor symmetry mixing turns on such a flux. Equivalently, due to the functional forms of ${\cal H}$ and $\Pi_{i}$ given in (89) and (91), we can think of this as a shift in $\Omega$ given by
@@
With the shifted versions of $v_{i}$ and $\Omega$ we can set $v_{i}=0$ and compute ${\cal Z}_{S^{3}}(\mu_{i})$ to determine a theories superconformal R-charge. However, as mentioned previously the flavor and R-charges of operators generically must be integral. This implies that the $\mu_{i}$ are also generically integral. Only when $g-1=0\ \text{mod}\ p$ can we choose a representative $v_{R}$ such that $\mathfrak{n}_{R}=0$; allowing R-charges and thus the $\mu_{\alpha}$ to be arbitrary real numbers. For the case of $S^{3}\simeq{\cal M}_{0,1}$ such a representative of the R-symmetry background exists. Specifically, we must set $(v_{R},\mathfrak{n}_{R})=(1,0)$ in order to perform F-maximization calculations, otherwise ${\cal Z}_{S^{3}}(\mu_{i})$ is ill-defined for arbitrary real $\mu_{i}$. Comparing our formulas (28) and (94), we seem to have two very different descriptions of the 3-sphere partition function which we could use for F-maximization. Fortunately, for our purposes they are equivalent
@@
In this work, when we deform the UV theory ${\cal T}[K,\emptyset]$ by monopole operators to arrive at ${\cal T}[K,\{{\cal O}_{I}\}]$, the effect can be understood as follows. Given a relevant, dressed, monopole operator which we can add to the superpotential, we must insist that its R-charge remains 2 after mixing. In most of the cases studied in this paper, the theories with gauge group $U(1)^{r}$ were deformed by $r-1$ monopoles. This gives $r-1$ constraints on the $r$ variables $\mu_{i}$ and thus there is a one parameter family of $\mu_{i}(\tilde{\nu})$ along which we perform F-maximization:
@@
Where $a_{i}$ will always be the normalized charge vector of the unbroken $U(1)_{A}$ subgroup and $c_{i}$ are constants determined by requiring the superpotential have $R=2$. In this sense determining the value $\vec{\mu}_{0}$ at which the theory will be superconformal is reduced to a one-dimensional extremization problem on $|{\cal Z}_{{\cal M}_{0,1}}(\tilde{\nu})|$. If we define the extremizing value to be $\nu_{*}$, then $\vec{\mu}_{0}\>=\vec{\mu}(\nu_{*})$ which is the reference point defining the affine line (21).

### v2

@@ 4.2 Unitary TFTs
[…]
which implies that the only local operator in a unitary TFT is the identity operator. Similarly to the non-unitary case, the UV gauge theory has a flavor symmetry whose rank is equal to the dimension of $\mathfrak{M}[K,\{{\cal O}_{I}\}]$. For some examples, the index is $1$ for all $\vec{\mu}\in\mathfrak{M}[K,\{{\cal O}_{I}\}]$ even though its dimension is non-zero. In the case, it is natural to expect that the UV flavor symmetry decouples in the IR. As a consequence, the superconformal index and the three-sphere free energy $F(\vec{\mu})$ defined in (32) do not depend on the mixing parameter $\vec{\mu}\in\mathfrak{M}[K,\{{\cal O}_{I}\}]$.
For a UV supersymmetric Wilson loop $W_{\vec{Q}}$ to be a simple object in the IR TFT, it must satisfy the conditions similar to those in (38):
$\begin{split}&i)\;\;\langle W^{+}_{\vec{Q}}\rangle_{\rm sci}(q,\vec{\mu},\vec{\zeta}=1)=0\textrm{ or }\pm q^{\mathbb{Z}/2}\;,\\ &ii)\;\;\langle W^{+}_{\vec{Q}}W^{-}_{\vec{Q}}\rangle_{\rm sci}(q,\vec{\mu},\vec{\zeta}=1)=1\;,\end{split}$ (68)
@@
If ${\cal T}[K,\{{\cal O}_{I}\}]$ with a positive definite $K$ has a mass gap and a UV loop $W_{\vec{Q}}$ flows to a simple object in the IR TFT, the half-index
@@ 4.2.2 $r=2$
[…]
The first line follows from the basic mirror symmetry reviewed in Section 51. Integrating out the two chiral fields, we obtain the pure CS theory in the infrared. As discussed in Dimofte:2017tpi , the deformed Dirichlet boundary condition for $\Phi_{1}$ breaks the supersymmetry spontaneously unless its boundary value is at the critical point of the superpotential.
@@
The half-index computation suggests that the only boundary operator that survives is the identity operator. It is interesting to note that we also have  [^14] It is tempting to conjecture that the deformed Dirichlet boundary condition flows to that corresponds to “closing the puncture” of $D^{2}\times S^{1}$ in the IR, which produces the partition function on $S^{2}\times S^{1}$.
@@ 4.3 Comparision with Zagier’s result
Let us now compare our lists with the result in Zagier:2007knq . The half-index $I_{A}[W_{\vec{Q}}]$ in (50) or $I_{\rm half}^{W_{\vec{Q}}}(q,\vec{\mu})$ in (69) corresponds to the modular Nahm sum formula $\chi_{(A,B,C)}$ in (3), with the identification
@@
provided that $B\in\mathbb{Z}^{r}$. We find that all the triples $(A,B,C)$ in the Zagier’s list, summarized in Table 1 (with $B\in\mathbb{Z}^{r}$), appear in our classification except for two cases.
@@
The ${\cal T}[K,\emptyset]$ has only one primitive 1/2 BPS monpole operator $V_{(0,1,1)}$. After the superpotential deformation, the gauge theory ${\cal T}[K,V_{(0,1,1)}]$ flows to an ${\cal N}=4$ SCFTs in the IR. The UV gauge theory has two $U(1)$ topological symmetries but the computation of superconformal index strongly suggests that one linear combintations of them acts trivially in the IR. If this happens, the example does not show up in our search due to our simplifying assumption (a) in section 3.2.
@@
Here $\alpha\in\mathbb{R}$ parametrize the mixing of R-symmetry with the decoupled $U(1)$ symmetry. One can check that the superconformal index is independent on $\alpha$. The half-index computation at $\alpha=0$ becomes
[…]
@@
[…]
@@
in Zagier’s list. The Nahm sum formula with (81) for both choices of $B$ gives
@@
This is a modular function but we do not know whether it can be identified with a vacuum character of any rational VOA. [^16] As coefficients of $\chi^{U(1)_{2}}_{1}(q)$ are all even, the coefficients of $q$-series are all integer-valued It can be reproduced from the half index of U3-1 with $a=2$ at $\vec{\mu}=\vec{0}$ decorated by a supersymmetric Wilson loop of charge $\vec{Q}=-(1,1,0)$. However, when computing the superconformal index:
@@ 5 Discussion
The gauge theory $T[K,\{\mathcal{O}_{I}\}_{I=1}^{r-1}]$ should be understood as a sequence of renormalization group (RG) flows:
$T[K,\emptyset]$ $\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{1}\;}T[K,\{\mathcal{O}_{1}\}]\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{2}\;}\ldots\xrightarrow{\;\delta\mathcal{W}=\mathcal{O}_{r-1}\;}T[K,\{\mathcal{O}_{I}\}_{I=1}^{r-1}]\;.$ (84)
For the RG flows to make sense, the superpotential at each step must be relevant:
@@
Here, $R^{(I-1)}_{0}=R_{*}+\vec{\mu}^{(I-1)}_{0}\cdot\vec{T}$ represents the superconformal R-charge of the theory $T[K,\{\mathcal{O}_{A}\}_{A=1}^{I-1}]$, which can be determined by extremizing the $F(\vec{\mu})$ over $\vec{\mu}\in\mathfrak{M}[K,\{\mathcal{O}_{1},\ldots,\mathcal{O}_{I-1}\}]$. There should be a proper sequence of the 1/2 BPS operators, $\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{r-1}$, to ensure that the relevance conditions are satisfied.
Verifying the relevance of the superpotential is a challenging task, even numerically, and will address it in future work. Instead, we will indirectly demonstrate that certain examples in our classification may violate this condition by highlighting inconsistencies in the IR fixed point. In the cases of 3-14-(2) and 3-20, the half indices $I_{A}[W_{\vec{Q}}]$ contain a factor $\chi_{F}^{-1}$, which does not correspond to any RCFT character known in literature. Additionally, the A and B-twisted partition functions on $S^{2}\times S^{1}$ are not equal to $1$ for the 3-14-(2) case. In the 3-20 case, the half-index $I_{A}[W_{(1,1,0)}]$ is expressed as a sum of two characters (modulo a $\chi_{F}^{-1}$), even though $W_{(1,1,0)}\in W^{\rm sim}_{A}$. Notably, both cases involve a non-primitive monopole operator: $V_{(-2,0,4)}$ in 3-14-(2) and $V_{(-2,2,2)}$ in 3-20, which we suspect to be irrelevant.
@@
In section 3.3, we have scanned UV Wilson loops which flow to simple objects in the IR TFT. Half-indices with the Wilson loops well matched to the RCFT characters of the corresponding primaries. But there are some exceptional cases that we couldn’t find the UV Wilson loop corresponding to an IR simple object: $\chi^{SM(2,12)}_{(1,3)}$ of 3-22, and $\chi^{W_{3}(3,7)}_{(3,3,1)}$ of 3-24.
In 3-22, we couldn’t find the UV Wilson loop corresponding to $\chi^{SM(2,12)}_{(1,3)}$. Instead, we find that
@@
The relation indicates that the UV Wilson loop flows to a non-simple line in the IR, which can be written as a linear combination of two simple lines. Suppose that the deformed Dirichlet boundary condition in the UV gauge theory maps to a holomorphic boundary condition in the IR topological theory denoted by $\langle\text{hol}|$, so that the vacuum character of the boundary algebra can be written as
@@
$q^{2}I_{A}[W_{(-2,2,1)}]=q^{4}I_{A}[W_{(-4,3,1)}]=q^{-3/28}\chi^{W_{3}(3,7)}_{(3,3,1)},$ (91)
but they are not simple: $\langle W^{-}_{(-2,2,1)}W^{+}_{(-2,2,1)}\rangle_{\rm sci}=\langle W^{-}_{(-4,3,1)}W^{+}_{(-4,3,1)}\rangle_{\rm sci}=3$. This suggests that the Wilson loops can be expressed as linear combinations of three IR simple objects. Based on our computations, we propose the following expressions for the two Wilson loops in terms of IR simple objects:
@@
Note that there are two primaries with a conformal dimension of 3/7-3/7 in the $W_{3}$ minimal model and the corresponding bulk simple objects are denoted by $\mathbf{L}_{-\frac{3}{7}}$ and $\widetilde{\mathbf{L}}_{-\frac{3}{7}}$. They are related to each other by a charge conjugation and they share the same conformal character. The above identification is also consistent with $\langle W^{-}_{(-2,2,1)}W^{+}_{(-4,3,1)}\rangle_{\rm sci}=-q^{-2}$.
@@ A.1 General structure of partition functions on $\mathcal{M}_{g,p}$
The effective twisted superpotential $\mathcal{W}(u,v)$ is broken down into two pieces, one originating from the Chern-Simons terms of the theory and one originating from the one-loop contribution of the $\Phi_{i}$ [^18] Importantly this is all done in the $U(1)_{-\frac{1}{2}}$ quantization” which is arises due to the parity anomaly that exists in 3D gauge theories as discussed in Closset:2018ghr .:
@@
where the A subscript indicates we are utilizing the A-twist background on $S^{2}$ as described in Closset:2017zgf . The flavor flux operator $\Pi_{i}(u,v)^{\mathfrak{n}_{i}}$ can also be inserted to compute the partition function in the presence of a nontrivial flavor flux $\mathfrak{n}_{\alpha}$ threaded through the base space $\Sigma_{g}$. Lastly, the gauge flux operator, $\Pi_{a}(u,v)$ is a rational function in the $x_{a}$ which defines the Bethe vacua [^19] For non-abelian gauge group $G$ the action of the Weyl group of $G$ on ${\cal S}_{BE}$ must be considered. as solutions to the system of polynomial equations:
@@
$\mathcal{Z}_{\mathcal{M}_{g,p}}(v;v_{R},\mathfrak{n}_{\alpha})=\sum_{\hat{u}\in\mathcal{S}_{BE}(v;v_{R})}\mathcal{H}(\hat{u},v)^{g-1}\mathcal{F}(\hat{u},v)^{p}\prod_{i}\Pi_{i}(\hat{u},v)^{\mathfrak{n}_{i}}.$ (103)
@@ A.2.1 $U(1)_{R}$ Holonomy and Flux in the 3D A-model
Recall that for a field with R-charge $r$ the Dirac quantization condition is $r\mathfrak{n}_{R}\in\mathbb{Z}$ and thus the A-model background generically enforces $r\in\mathbb{Z}$. However, examining the conditions (105), we see that when $p\neq 0$, g1=0(modp)g-1=0\ ($\text${mod}\ p), and $v_{R}$ is integral [^20] One can find conditions for $v_{R}$ half-integral, namely $g-1=\frac{p}{2}\ (\text{mod}\ p)$, but this will not be important here., a specific representative $v_{R}$ can be chosen such that $\mathfrak{n}_{R}=0$ in which case we are free to allow $r\in\mathbb{R}$. This is crucial for F-maximization where the R-charge must be allowed to take on arbitrary real values.
@@ A.2.2 F-maximization in the 3D A-model
[…]
@@
With the shifted versions of $v_{i}$ and $\Omega$ we can set $v_{i}=0$ and compute ${\cal Z}_{S^{3}}(\mu_{i})$ to determine a theories superconformal R-charge. However, as mentioned previously the flavor and R-charges of operators generically must be integral. This implies that the $\mu_{i}$ are also generically integral. Only when g1=0modpg-1=0\ $\text${mod}\ p can we choose a representative $v_{R}$ such that $\mathfrak{n}_{R}=0$; allowing R-charges and thus the $\mu_{\alpha}$ to be arbitrary real numbers. For the case of $S^{3}\simeq{\cal M}_{0,1}$ such a representative of the R-symmetry background exists. Specifically, we must set $(v_{R},\mathfrak{n}_{R})=(1,0)$ in order to perform F-maximization calculations, otherwise ${\cal Z}_{S^{3}}(\mu_{i})$ is ill-defined for arbitrary real $\mu_{i}$. Comparing our formulas (33) and (103), we seem to have two very different descriptions of the 3-sphere partition function which we could use for F-maximization. Fortunately, for our purposes they are equivalent
@@
[…]
@@
[…]